OpenStudy (anonymous):

question:

OpenStudy (anonymous):

Let L: \[R ^{2} -> R ^{2} \] be the linear transformation defined by \[L[\left(\begin{matrix}a \\ b\end{matrix}\right)]= \left[\begin{matrix}1 & 2 \\ 2 & 4\end{matrix}\right] \left(\begin{matrix}a \\ b\end{matrix}\right)\] a)find ker L b) find a set of vectors spanning range L.

OpenStudy (anonymous):

just explain the procedure how to do it. @lalaly

OpenStudy (unklerhaukus):

what does ker mean

OpenStudy (anonymous):

i guess its just L . ker is a printing mistake sorry for that uncle

OpenStudy (unklerhaukus):

?

OpenStudy (turingtest):

kernel

OpenStudy (anonymous):

maybe i m sure about kernel its only written ker L

OpenStudy (anonymous):

i just need the way to solve it

OpenStudy (turingtest):

I'm looking over my notes, I always forget this stuff

OpenStudy (anonymous):

no problem

OpenStudy (turingtest):

oh, kernel is the same thing as null space for a linear transformation i.e. the set of all \(\vec x\) in \(A\vec x=\vec 0\) is the kernel, so just solve that matrix to find part a)

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

that is, solve\[ \left[\begin{matrix}1 & 2 \\ 2 & 4\end{matrix}\right] \left(\begin{matrix}a \\ b\end{matrix}\right)=0\]

OpenStudy (anonymous):

then we'll get two eqs right ???

OpenStudy (turingtest):

yes, from which we will get the two vectors that will make up the basis of our kernel

OpenStudy (turingtest):

at least I think that is what they want

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

about what about part B ???

OpenStudy (turingtest):

consulting notes again...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

hm... I'm perhaps a bit confused because my notes don't seem to agree with wikipedia. I think the dimension of the null space, i.e. the nullity is the kernal in that case, the answer to part a) would be the number of vectors that make up the basis for the null space of your transformation matrix the answer to part b) must then be the basis itself I guess, which is the set of vectors that span \(A\vec x=\vec0\)

OpenStudy (anonymous):

ok thank you i got it :) i just need some practice now

OpenStudy (turingtest):

yeah, me too :) welcome, good luck